for # 7 i dont get why com only works on haning steel ball..
so there is no c.o.k.e?

for #8 i dont know what this question is asking me. first how do i tell if this question is asking me for impulse? it doesnt say anything about to solve for impulse...

i get the part where finding areas, but how do i know this is asking for impulse and initial velocity?

what where that 6*10^3 came from,, so damn confused on #8
explain in baby step because i dont understand this problem

#9 i got it.. .
Answer #1

#7 The answer says that if : at the other end two ball pop with half the speed, kinetic energy would not be conserved.

say taking two balls of mass m each as system,

change in KE = 0.5*m*v^2 - 0.5*m * (0.5v)^2 - 0.5*m * (0.5v)^2 = 0.25 *m *v^2

but if we take the case of other end ball popping with the same speed,

change in KE = 0.5*m*v^2 +0 - 0+ 0.5*m*v^2 =0

#8 Actually the question is asking for initial velocity (i,e touch down speed) and the final vel =0 from the relation: dp/dt = F (change in momentum / time )

=> deltaP= F*t (F*t is the area of given graph as u can see the unit of area is N*s)

since Vf=0 , so deltaP= m*Vi = F*t= 6*10^5 Ns

   => 6*10^3kg *Vi =   6*10^5 Ns

=> Vi= 100 m/s

touch down speed of plane is the speed of plane when its main gear wheel touches the surface, that implies its initial velocity.

6*10^3 kg is the mass of the aircraft.( mass* Vi = initial momentum of aircraft)